OpenStudy (anonymous):

\left(\begin{matrix}3 \\ sqrt{2}\end{matrix}\right)

OpenStudy (anonymous):

\left(\begin{matrix}3 \\ sqrt{2}\end{matrix}\right)

OpenStudy (saifoo.khan):

o.O

OpenStudy (anonymous):

\[\left(\begin{matrix}3 \\ \sqrt{2}\end{matrix}\right)\]